10/05/2021



T5:  VPhot submissins are not working correctly (and on multiple systems. Under investigation).


T7: replacement camera was dead on arrival to Spain, due to rough handling. We are currently cannibalising this camera for parts for the original camera. Camera repair completed. Pending star tests.


T8; Original FLI Microline 16803 mono camera is being refitted. Calibration, configuration and star testing to do.


T9: The Microline 16803 camera has a faulty usb port. To be sent to FLI USA for repairs. A temporary FLI ML8300C Colour camera is being fitted to T9.


T11: Calibration frames being redone.


T24; The camera CCD window has developed a 'haze' (not condensation) and has been sent for cleaning. Problems with clean room availability. Cleaning delayed.


T30  is requiring mount power cycling to clear slew errors. USB to serial converter replaced. Under observation.


T31 mount has an encoder error that is not allowing slews or homing. Under investigation.


T33 has been removed from SSO and is no longer available on the network.


T69; all components have finally been received; assembly in progress